please help, I don't know how to do this: find the value of $2500 invested at 4.5% annual interest for 10 years if interest is compounded quarterly, continuously
Answer by rfer(16322)

A=P(1+r/n)^nt
A=2500(1+.045/4)^4*10
A=2500(1.01125)^40
A=2500(1.5643769)
A=$3910.94
-----------------
A=Pe^rt
A=2500(2.7183)^0.045*10
A=2500(1.5832277)
A=$3958.07